find the slope of the line through the points(-14,-19), (-4,14)

Answer by stanbon(75887)

find the slope of the line through the points(-14,-19), (-4,14)
--------
slope = (14--19)/(-4--14) = (14+19)/(-4+14) = 33/10
===============
Cheers,
Stan H.
===============